Suppose a retailer accepts four types of credit cards and lists the types used by the last 50 customers. Construct a frequency distribution for each type of credit card.

Discover Visa Visa Am. Express Visa
Visa Visa Am. Express MasterCard Visa
Am. Express MasterCard Visa Visa Visa
Visa Am. Express Am. Express MasterCard Visa
MasterCard Visa Discover Am. Express Discover
Visa Am. Express Discover Visa MasterCard
Visa Visa Visa Visa MasterCard
MasterCard Am. Express Visa MasterCard Visa
MasterCard Discover MasterCard Visa Visa
MasterCard Discover Am. Express Discover Visa

SOLUTION:
Type Frequency
MasterCard
Visa
American Express
Discover
1
Expert's answer
2019-09-10T10:58:42-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \hdashline Type & Frequency \\ \hdashline
 Am.Express & 9 \\
 \hdashline
 Discover & 7 
\\ \hdashline
 MasterCard & 11 \\
 \hdashline
 Visa & 23 
\end{array}"

Check


"9+7+11+23=50"

The proportion of observations in a category is called the relative frequency of the category.


"Relative \ \ Frequency={Frequency \over Sum\ of\ all\ frequencies}"

Compute the relative frequency for each type of credit card


"\def\arraystretch{1.5}
 \begin{array}{c:c: c:}
 \hdashline Type & Frequency & Relative\ Frequency \\ \hdashline
 Am.Express & 9 & 9/50=0.18\\
 \hdashline
 Discover & 7 & 7/50=0.14 
\\ \hdashline
 MasterCard & 11 & 11/50=0.22 \\
 \hdashline
 Visa & 23 & 23/50=0.46
\end{array}"

Check


"0.18+0.14+0.22+0.46=1"